### v1

# Adaptive Cache Enhancement for Test-Time Adaptation of Vision-Language Models

Abstract
Vision-language models (VLMs) exhibit remarkable zero-shot generalization but suffer performance degradation under distribution shifts in downstream tasks, particularly in the absence of labeled data. Test-Time Adaptation (TTA) addresses this challenge by enabling online optimization of VLMs during inference, eliminating the need for annotated data. Cache-based TTA methods exploit historical knowledge by maintaining a dynamic memory cache of low-entropy or high-confidence samples, promoting efficient adaptation to out-of-distribution data. Nevertheless, these methods face two critical challenges: (1) unreliable confidence metrics under significant distribution shifts, resulting in error accumulation within the cache and degraded adaptation performance; and (2) rigid decision boundaries that fail to accommodate substantial distributional variations, leading to suboptimal predictions. To overcome these limitations, we introduce the Adaptive Cache Enhancement (ACE) framework, which constructs a robust cache by selectively storing high-confidence or low-entropy image embeddings per class, guided by dynamic, class-specific thresholds initialized from zero-shot statistics and iteratively refined using an exponential moving average and exploration-augmented updates. This approach enables adaptive, class-wise decision boundaries, ensuring robust and accurate predictions across diverse visual distributions. Extensive experiments on 15 diverse benchmark datasets demonstrate that ACE achieves state-of-the-art performance, delivering superior robustness and generalization compared to existing TTA methods in challenging out-of-distribution scenarios.
Introduction
Vision-language models (VLMs) such as CLIP (Radford et al. 2021a) and ALIGN (Jia et al. 2021) exhibit remarkable zero-shot abilities in a wide range of tasks, including classification, retrieval, and segmentation (Zhou et al. 2021; Khattak et al. 2022; Xie et al. 2023; Iscen et al. 2024; Jiahaoli et al. 2024; Zhu et al. 2023). These models are trained to learn the semantic relationship between visual and textual information and can be used for image-text matching during inference. However, their effectiveness often decreases when handling unlabeled test data from domains that substantially differ from their training distribution (Agarwal et al. 2021; Menon, Chandratreya, and Vondrick 2024; Parashar et al. 2024).
Figure 1: Noisy CLIP predictions on different views of “dog” image. CLIP can be either overconfident on simple view but fail on strong augmentation.
To address the challenge of performance decline of VLMs to distribution shifts, test-time adaptation (TTA) has emerged as a strong and effective strategy for improving their adaptability in out-of-distribution (OOD) cases. Unlike supervised fine-tuning or prompt learning, which requires expensive labeled data, TTA exploits unlabeled test data to dynamically adjust model predictions during inference. This characteristic makes TTA suitable for real-world applications, such as autonomous systems and medical imaging, where annotated datasets are often limited.
Prompt-based TTA methods (Shu et al. 2022a; Feng et al. 2023a; Hassan et al. 2023) reduce the prediction entropy of VLMs by optimizing domain-specific prompts. However, their prompting optimization requires repeated backpropagation through the encoders of VLMs, resulting in substantial computational cost and hence limiting their scalability in resource-limited settings. Cache-based TTA strategies (Karmanov et al. 2024a; Zhang et al. 2024b, c, a) dynamically assemble a cache of high-confidence samples guided by VLM prediction entropy metrics, thus fine-tuning predictions with minimal parameter changes. For example, TDA (Karmanov et al. 2024a) uses adaptive caching to enhance the calibration of predictions, while DMN (Zhang et al. 2024c) utilizes short- and long-term memory banks to leverage past insights. Similarly, DPE (Zhang et al. 2024a) improves multimodal alignment through dual-modal residual learning.
However, as demonstrated in Figure 1, CLIP predictions vary significantly under different augmentations, making the entropy-based sample selection unreliable and potentially hindering the adaptability of the model (Khan and Fu 2024; Farina et al. 2024). Furthermore, the decision boundaries of the VLMs in target domains become inconsistent. Even with TTA, these models frequently fail to adjust the decision boundaries to accommodate domain shifts, resulting in fluctuating predictions. Although several methods (Zhang et al. 2024c, a) attempt to overcome this issue by using mean textual prototypes from enriched prompts, they are limited in handling the variability from intra-class diversity shifts.
To address the above challenges, we propose the Adaptive Cache Enhancement (ACE) framework, which constructs a robust cache by selectively storing high-confidence or low-entropy image embeddings for each class, directed by dynamic, class-specific thresholds initialized from zero-shot statistics and iteratively refined with an exponential moving average and exploration-augmented updates. By excluding over-confident samples from the cache, our method avoids overfitting, allows for adaptable class-specific decision boundaries, and ensures reliable and precise predictions across a broad range of data distributions. To this end, we make the following contributions in our work.
  • We propose ACE, a novel framework for boosting up the performance of VLMs in domain shifts.
  • We introduce an adaptive and class-wise thresholding mechanism to enhance the reliability of predictions across diverse data distributions. Our thresholding mechanism is initialized from zero-shot statistics and refined via exponential moving average and exploration-augmented updates.
  • We validate the robustness and generality of our method through extensive experiments on 15 benchmark datasets and show the superiority of our ACE over current research in challenging OOD scenarios.
Related Work
Vision-Language Models (VLMs)
Large-scale pre-trained VLMs such as CLIP (Radford et al. 2021a) have demonstrated substantial proficiency in acquiring transferable representations that span both imagery and text. Extensive research has been conducted to optimize these models for downstream applications. Initial efforts, including prompt learning strategies (Zhou et al. 2021, 2022a), achieve notable success by implementing minimal supervision in few-shot learning environments. Recent methods, e.g., Tip-Adapter (Zhang et al. 2021) and TaskRes (Yu et al. 2023), have effectively utilized feature memories derived from a limited number of labeled instances for rapid adaptation. These memory-based approaches have gained popularity due to their ability to facilitate quick domain adaptation with minimal computational burden, eliminating the requirement for full model fine-tuning. Nevertheless, their dependency on labeled target-domain data limits their applicability in real-world contexts, where such data are often not available. Our research focuses on test-time adaptation, which compels VLMs to acclimate to new domains during the inference stage without access to labeled data. This scenario closely reflects the real-world setting, advancing the development of flexible vision-language systems.
Test-Time Adaptation (TTA) for VMLs
TTA enhances the adaptability of VLMs to unfamiliar target domains by exclusively leveraging unlabeled test data (Shu et al. 2022a; Ma et al. 2023; Farina et al. 2024; Zhao et al. 2024). Current methods are generally divided into prompt-based and cache-based categories. Prompt-based techniques refine textual embeddings throughout inference. For example, TPT (Shu et al. 2022a) refines prompts to maintain consistency across multiple augmented image views. DiffTPT (Feng et al. 2023a) applies diffusion-driven augmentation techniques to enhance view diversity, thus improving the adaptability of VLMs to unfamiliar test data. Conversely, cache-based methods utilize historical, instance-specific data extracted from test samples. Specifically, TDA (Karmanov et al. 2024a) creates caches containing positive and negative features to improve CLIP’s predictions through aggregated similarity measurements. DMN (Zhang et al. 2024c) combines static and dynamic memory networks, employing an advanced interaction mechanism to effectively leverage both few-shot and accumulated unlabeled data for robust adaptation. More recently, DPE (Zhang et al. 2024a) refines visual and textual prototypes in parallel to capture domain-specific semantics more precisely. Although prompt-based techniques are appreciated for their simplicity, cache-based strategies often achieve superior adaptability and computational efficiency by modeling target domain distributions through lightweight, memory-based mechanisms at test time. Expanding on these cache-based methods, our research specifically aims at improving cache quality through reliable sample selection and enhancing predictive strength by adaptively adjusting decision boundaries.
Figure 2: Overview of the ACE Method. We introduce the Class-wise Adaptive Threshold Enhancement module, which refines the threshold online to allow more correct samples but low-confidence samples to be cached. By simultaneously minimizing entropy loss for the prototype residuals, and creating highly reliable caches, ACE enhances the overall multimodal generalization and robustness.
Proposed Method
We propose an effective and reliable TTA strategy that includes a cache data structure, prototype residual learning modules, a flexible class-wise thresholding mechanism, and an adaptive threshold refinement module based on the statistic of pseudo-labels. We realize our method with CLIP (Radford et al. 2021a), a state-of-the-art VLM model, and DPE (Zhang et al. 2024a), a prototype residual vector learning technique.
Preliminaries
Zero-shot CLIP Prediction.
CLIP is pre-trained on a large collection of image-text pairs to align visual and textual information within an embedding space. CLIP can be trained by optimizing a contrastive loss that maximizes the cosine similarity between matching image-text pairs. CLIP comprises a visual encoder $\mathcal{E}_{v}$ and a text encoder $\mathcal{E}_{t}$, which map images and text into a shared latent space $\mathbb{R}^{d}$. At inference time, an image $x$ is classified by measuring the similarity between its visual representation $\mathbf{z}=\mathcal{E}_{v}(x)$ and textual embeddings $\textbf{t}^{c}=\mathcal{E}_{t}(\mathcal{P}^{c})$, where $\mathcal{P}^{c}$ is a text prompt describing a class $c$, and $C$ is the total number of classes. The zero-shot CLIP prediction probability is then given by:
$p^{c}_{\text{CLIP}}=\frac{\exp\left(\mathbf{z}\cdot\textbf{t}^{c}/\tau\right)}{\sum_{j=1}^{C}\exp\left(\mathbf{z}\cdot\textbf{t}^{j}/\tau\right)}$ (1)
where $\tau$ is the temperature parameter, typically set to 0.01.
Cache-based TTA.
Inspired by the key-value storage framework introduced in TIP-Adapter (Zhang et al. 2022), several cache-based TTA approaches (Karmanov et al. 2024a; Zhang et al. 2024a, b) have been proposed, which dynamically store and update high-confidence feature-label pairs during inference. The cache is organized as a priority queue, where the priority of each entry is determined by the entropy of CLIP’s predictions.
Let $\mathbf{W}_{\mathrm{CLIP}}=\left[\textbf{t}^{1},\textbf{t}^{2},\ldots,\textbf{t}^{C}\right]^{\top}$ be a stack of text embeddings for all classes. Each cache item is represented as:
$\left\{\mathbf{F}_{\text{cache}},\mathbf{L}_{p},\mathcal{H}\left(\mathbf{F}_{\text{cache}}\mathbf{W}_{\text{CLIP}}^{\top}\right)\right\},$ (2)
[…]
$\mathcal{H}(p)=\left(-\sum_{i=1}^{C}p_{i}\log p_{i}\right)$ (3)
As test samples arrive sequentially, new entries are inserted into cache slots corresponding to their pseudo-labels. When the cache exceeds its capacity, entries with the highest entropy are evicted. During inference, a test feature $\mathbf{z}$ retrieves related cache items, and the final prediction is calculated by combining the CLIP logits with those derived from the cache:
$\text{logits}_{\mathrm{cls}}(\mathbf{z})=\mathbf{z}\mathbf{W}_{\text{CLIP}}^{\top}+\mathcal{F}(\mathbf{z}\mathbf{F}_{\text{cache}}^{\top})\mathbf{L}_{p},$ (4)
where $\mathcal{F}(x)=\alpha\exp(-\beta(1-x))$ is a modulation function with a scaling factor $\alpha$ and a sharpness parameter $\beta$.
Following the DPE approach Zhang et al. (2024a), the feature for each cache slot, $\mathbf{F}_{\text{cache}}$, is computed as the mean of all stored features for that class, forming increasingly compact and representative class-specific prototypes as more data accumulates. The cache item is then constructed as:
$\left\{\mathbf{F}_{\text{cache}},\mathbf{L}_{p},\mathcal{H}\left(\mathbf{F}_{\text{cache}}\mathbf{W}_{\text{CLIP}}^{\top}\right)\leq\mathcal{T}\right\},$ (5)
where $\mathcal{T}$ is an entropy threshold for the caching decision.
Adaptive Cache Enhancement (ACE)
Cache-based strategies demonstrate notable effectiveness in various domains. However, they are grappling with two substantial issues. First, entropy is not a reliable indicator when the data distribution changes. Second, fixed decision boundaries are not adaptive to distributional changes. To address these challenges, we introduce two specialized modules that specifically address each issue while synergistically strengthening each other as illustrated in Figure 2.
Curriculum Thresholding
The task of dynamically setting class-specific thresholds in alignment with the evolving learning progress of the model is challenging. Ideally, the threshold for each class can be set specifically based on evaluation metrics, e.g., recognition accuracy or entropy as,
$\mathcal{T}_{t}(c)=m_{t}(c)\cdot\mathcal{T}_{0}(c),$ (6)
where $\mathcal{T}_{t}(c)$ is the adaptive threshold for class $c$ at time step $t$, and $m_{t}(c)$ denotes the corresponding metric, whether it is accuracy or entropy. This approach ensures that classes with poorer performance, indicated by lower accuracy rates, are designated smaller thresholds (or conversely, higher entropy leads to larger thresholds), thereby permitting a greater number of their instances to be retained in the cache during the testing phase. However, dynamically adjusting these thresholds throughout the training process requires continuous accuracy estimates at time increments $t$, which impedes training efficiency.
In this study, we introduce Curriculum Thresholding, a thresholding mechanism that evaluates the learning status of classes without introducing any additional inference steps and/or requiring extra annotations. Central to our hypothesis is the postulation that when subjected to a high threshold, the quantity of confidently classified predictions ascribed to a particular class serves as a robust indicator of the learning progress for that class. Specifically, we proceed with the following steps:
$i\noindent\sigma_{t}(c)$ $=\sum_{n=1}^{N}\mathds{1}\left(\max\left(p(y|x)\right)\geq\mathcal{T}_{t-1}(c)\right)\cdot\mathds{1}\left(\arg\max(p(y|x))=c\right),$
or $=\sum_{n=1}^{N}\mathds{1}\left(\min\left(e(y|x)\right)\leq\mathcal{T}_{t-1}(c)\right)\cdot\mathds{1}\left(\arg\max\left(p(y|x)\right)=c\right),$
where $\sigma_{t}(c)$ measures the number of samples confidently predicted as class $c$ at time $t$, and $p(y|x)$ and $e(y|x)$ are the probability and entropy of the model to sample $x$.
We then normalize $\sigma_{t}(c)$ as,
$\sigma_{t}(c)\leftarrow\frac{\sigma_{t}(c)}{\max_{c}\sigma_{t}}$ (7)
Intuitively, a higher $\sigma_{t}(c)$ indicates that the model has learned class $c$ more effectively. Next, we compute an adaptive threshold $\mathcal{T}_{t}(c)$ for class $c$ at time step $t$ using the Exponential Moving Average (EMA) as,
$\mathcal{T}_{t}(c)=\delta\mathcal{T}_{t-1}(c)+(1-\delta)\cdot m_{t}(c)$ (8)
where $m_{t}(c)=\sigma_{t}(c)\cdot m_{t-1}(c)$ is the latest performance metric (accuracy or entropy) and $\delta$ is an EMA factor.
Nevertheless, considering that Eq. (8) is executed solely during the testing phase, it should be anticipated that initial performance might not be optimal. To improve the adaptability and reduce overfitting of the model, particularly for classes well-represented in the dataset, the initial threshold at time $t=0$ is determined using statistics from zero-shot learning. This statistic is simply the average of the entropy and probability across all samples in the test set. This initialization effectively captures the true behavior of both the model and the test set, ensuring reliable and meaningful results. It is important to highlight that incorporating CT is essentially cost-effective. In practice, whenever the confidence level of a sample’s prediction surpasses the threshold $\mathcal{T}$, the sample, along with its predicted class, is flagged. Flagged instances will be retained in the cache in subsequent iterations.
Online Thresholding Adaptation
The CLIP model exhibits limitations in terms of its applicability across all datasets. Consequently, in scenarios such as zero-shot evaluations or even during TTA, predictions may not always be made for certain classes. It is critical to monitor classes that have been encountered infrequently or not at all. For these infrequently seen or never-seen classes, we implement a strategy for more intensive threshold exploration. This approach enhances the diversity of the cache and accelerates the adaptation to classes that exhibit a long-tailed distribution. To streamline our analysis, we categorize any class with fewer than 10 predictions as rarely seen classes.
$\text{Never seen}:\mathcal{T}_{t}(c)$ $=\sum_{1}^{C}\mathds{1}\left(M_{c,t}=0\right)*\mathcal{T}_{t}(c)*(1-\gamma)$
$\text{Rarely seen}:\mathcal{T}_{t}(c)$ $=\sum_{1}^{C}\mathds{1}\left(M_{c,t}\leq 10\right)*\mathcal{T}_{t}(c)*(1-\gamma*0.5)$
where $\gamma$ is an adaptation rate and $M_{c,t}$ is the number of samples of class $c$ at time $t$ in the cache.
The online thresholding adaption ensures that the best-learned class reaches the full threshold $\mathcal{T}$, while the less-learned classes get proportionally lower thresholds, encouraging more of their samples to be trained. As the training progresses, the model improves and the thresholds converge towards $\mathcal{T}$. Notably, the thresholds are not guaranteed to increase over time; they may decrease if predictions become less confident in later iterations.
Method ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Average OOD Average
CLIP-ResNet-50 (Radford et al. 2021b) 58.16 21.83 51.41 56.15 33.37 44.18 40.69
Ensemble 59.81 23.24 52.91 60.72 35.48 46.43 43.09
CoOp (Zhou et al. 2022b) 63.33 23.06 55.40 56.60 34.67 46.61 42.43
TPT (Shu et al. 2022b) 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT (Feng et al. 2023b) 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA (Karmanov et al. 2024b) 61.35 30.29 55.54 62.58 38.12 49.58 46.63
TPS (Sui, Wang, and Yeung-Levy 2024) 61.47 30.48 54.96 62.87 37.14 49.38 46.36
DMN-ZS (Zhang et al. 2024d) 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE (Zhang et al. 2024a) 63.41 30.15 56.72 63.72 40.03 50.81 47.66
CRG (Zhai et al. 2025) 65.26 29.69 56.07 - - - -
ACE-Probability 63.75 30.09 56.84 63.95 40.15 50.94 47.76
ACE-Entropy 63.70 30.20 56.89 63.82 40.25 50.97 47.79
CLIP-ViT-B/16 (Radford et al. 2021b) 66.73 47.87 60.86 73.98 46.09 59.11 57.20
Ensemble 68.34 49.89 61.88 77.65 48.24 61.20 59.42
CoOp (Zhou et al. 2022b) 71.51 49.71 64.20 75.21 47.99 61.72 59.28
TPT (Shu et al. 2022b) 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT (Feng et al. 2023b) 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA (Karmanov et al. 2024b) 69.51 60.11 64.67 80.24 50.54 65.01 63.89
TPS (Sui, Wang, and Yeung-Levy 2024) 70.19 60.08 64.73 80.27 49.95 65.04 63.76
DMN-ZS (Zhang et al. 2024d) 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE (Zhang et al. 2024a) 71.91 59.63 65.44 80.40 52.26 65.93 64.43
CRG (Zhai et al. 2025) 75.01 63.67 64.66 - - - -
ACE-Probability 71.93 59.71 65.46 80.68 52.60 66.08 64.61
ACE-Entropy 71.93 59.62 65.51 80.72 52.66 66.09 64.63
Table 1: Performance comparisons on robustness to natural distribution shifts. We present top-1 accuracy (%) results for all evaluated methods employing both ResNet-50 and ViT-B/16 visual backbones of CLIP. The best and second-best results are highlighted in bold and underlined, respectively.
Test-Time Logit Inference and Loss Function.
We adopt DPE, the residual learning method proposed by Zhang et al. (2024d), to define textual and visual embedding prototypes as $\mathbf{t}_{c}=\frac{1}{S}\sum_{i=1}^{S}\mathcal{E}_{t}(\mathcal{P}^{c}_{i})$ and $\mathbf{v}_{c}=\frac{1}{S_{c}}\sum_{m=1}^{S_{c}}\mathbf{z}^{c}_{m}$. Here, $S$ is the number of prompts, $S_{c}\leq M$ is the total count of image features stored in the queue corresponding to class $c$, and $\mathbf{z}$ is the image features of $x$ that surpass the threshold $\mathcal{T}$. Once the two sets of multi-modal prototypes have been adjusted based on the last test sample, they are used to serve as the initial configuration for further refinement with the current test input. The preliminary outcomes derived from these steps are subsequently utilized to fine-tune the prototypes, making them adapted to $x$ as,
$\mathbf{t}_{c}\leftarrow\frac{\mathbf{t}_{c}+\hat{\mathbf{t}}_{c}}{\|\mathbf{t}_{c}+\hat{\mathbf{t}}_{c}\|},\quad\mathbf{v}_{c}\leftarrow\frac{\mathbf{v}_{c}+\hat{\mathbf{v}}_{c}}{\|\mathbf{v}_{c}+\hat{\mathbf{v}}_{c}\|}.$ (9)
where $\hat{\mathbf{t}}_{c}\in\mathbb{R}^{C\times d}$ and $\hat{\mathbf{v}}_{c}\in\mathbb{R}^{C\times d}$ are the learnable residual parameters that are initialized as zero and are used to optimize the prototypes for each given test input $x$. Similarly to Zhang et al. (2024d), we optimize these residual parameters to promote consistent predictions in a total of $N$ different augmented views of each given test image $x$ using the objective of unsupervised entropy minimization,
$\mathcal{L}_{\mathtt{aug}}$ $=\mathcal{H}(\mathbb{P}_{\mathtt{ACE}}(x))$
$=-\sum_{c=1}^{C}\mathbb{P}_{\mathtt{ACE}}(y=y_{c}|x)\log\mathbb{P}_{\mathtt{ACE}}(y=y_{c}|x)$ (10)
where
$\mathbb{P}_{\mathtt{ACE}}(x)$ $=\frac{1}{\rho N}\sum_{n=1}^{N}\mathds{1}[\mathcal{H}\left(\mathbb{P}_{\mathtt{Proto}}(\mathcal{A}_{n}(x)\right)\leq\mathcal{T}_{e}]\,\mathbb{P}_{\mathtt{Proto}}(\mathcal{A}_{n}(x))$
or $=\frac{1}{\rho N}\sum_{n=1}^{N}\mathds{1}[p\left(\mathbb{P}_{\mathtt{Proto}}(\mathcal{A}_{n}(x)\right)\geq\mathcal{T}_{p}]\,\mathbb{P}_{\mathtt{Proto}}(\mathcal{A}_{n}(x))$
where $\mathcal{A}_{n}(x)$ is an augmented view of $x$.
We adopt the alignment loss in DPE,
$\mathcal{L}_{\mathtt{align}}=\frac{1}{C}\sum_{c=1}^{C}-\log\left(\frac{\exp(\textbf{t}_{c}^{\,\,\top}\textbf{v}_{c})}{\sum\nolimits_{c^{\prime}}\exp(\textbf{t}_{c}^{\,\,\top}\textbf{v}_{c^{\prime}})}\frac{\exp(\textbf{t}_{c}^{\,\,\top}\textbf{v}_{c})}{\sum\nolimits_{c^{\prime}}\exp(\textbf{t}_{c^{\prime}}^{\,\,\top}\textbf{v}_{c})}\right)$ (11)
The final objective for optimizing the multi-modal prototypes $\mathbf{t},\mathbf{v}$ is defined as,
$\hat{\mathbf{t}},\hat{\mathbf{v}}=\arg\min_{\mathbf{t,v}}\{\mathcal{L}_{\mathtt{aug}}+\lambda\mathcal{L}_{\mathtt{align}}\},$ (12)
After optimizing the prototypes for each test sample $x$, we evolve the online textual prototypes $\mathbf{t}$, and update the priority queues to re-compute the visual prototypes $\mathbf{v}$. Then, we perform the testing as usual as in Eq. (4).
Method Aircraft Caltech Cars DTD EuroSAT Flower Food101 Pets SUN397 UCF101 Average
CLIP-ResNet-50 15.66 85.88 55.70 40.37 23.69 61.75 73.97 83.57 58.80 58.84 55.82
CoOp (Zhou et al. 2021) 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 56.18
TPT (Shu et al. 2022a) 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 57.66
DiffTPT (Feng et al. 2023c) 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 59.85
TDA (Karmanov et al. 2024a) 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.03
DPE (Zhang et al. 2024a) 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
CRG (Zhai et al. 2025) 18.09 90.12 57.92 51.89 46.80 71.09 75.76 85.91 63.11 63.58 62.42
ACE-Probability 21.33 90.91 59.84 55.20 51.02 70.28 76.83 87.95 65.46 61.83 64.07
ACE-Entropy 21.48 90.99 60.00 53.90 48.54 70.08 76.91 87.54 65.38 62.60 63.74
CLIP-ViT-B/16 23.67 93.35 65.48 44.27 42.01 67.44 83.65 88.25 62.59 65.13 63.58
CoOp (Zhou et al. 2021) 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88
TPT (Shu et al. 2022a) 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT (Feng et al. 2023c) 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47
TDA (Karmanov et al. 2024a) 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
DPE (Zhang et al. 2024a) 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
CRG (Zhai et al. 2025) 26.58 93.57 66.89 56.38 59.81 75.94 85.95 91.20 68.36 70.31 69.50
ACE-Probability 29.46 95.13 68.72 56.38 54.28 77.47 85.88 91.47 70.83 73.62 70.32
ACE-Entropy 29.49 95.33 68.55 56.56 53.63 77.30 85.88 91.47 70.88 73.72 70.28
Table 2: Performance comparisons on cross-dataset generalization across 10 fine-grained datasets. The best and second-best results are highlighted in bold and underlined, respectively.
Experiments
In this section, we assess the robustness of the proposed ACE method to natural distribution shifts and its cross-dataset generalization across 15 different datasets. In addition, we compare the test-time efficiency of ACE against existing approaches. Lastly, we conduct ablation studies to examine the impact of various components and design choices.
Experimental Setup
Datasets
Following prior research (Karmanov et al. 2024a; Zhang et al. 2024a), we evaluate our method in two main benchmarking scenarios: cross-dataset generalization and robustness to natural distribution shifts. (1) For the cross-dataset generalization scenario, we carry out evaluations across 10 varied recognition datasets, including FGVC Aircraft (Maji et al. 2013a), Caltech101 (Fei-Fei, Fergus, and Perona 2004), Stanford Cars (Krause et al. 2013a), DTD (Cimpoi et al. 2013), EuroSAT (Helber et al. 2019), Flowers102 (Nilsback and Zisserman 2008a), Food101 (Bossard, Guillaumin, and Van Gool 2014), Oxford Pets (Parkhi et al. 2012a), SUN397 (Xiao et al. 2010a), and UCF101 (Soomro 2012). These datasets collectively offer a comprehensive benchmark to assess the adaptability and generality of models across different datasets. (2) To evaluate the robustness of our method against natural distribution shifts, we test our method with ImageNet (Deng et al. 2009) and its out-of-distribution variants, including ImageNet-A (Hendrycks et al. 2021b) and ImageNet-V2 (Recht et al. 2019). This part of the evaluation assesses our method’s robustness under varying distribution shifts. For datasets with particularly large test sets, e.g., Food101, ImageNet, SUN397, ImageNet-V2, and ImageNet-A, we sample 2,000 test instances for evaluation, adhering to the experimental protocol in DiffTPT (Feng et al. 2023c).
Implementation details
Following previous studies (Karmanov et al. 2024a; Zhang et al. 2024a), we select ResNet-50 and ViT-B/16 as visual encoders for CLIP, employing hand-crafted prompts as described in the supplementary material. In line with the method by (Shu et al. 2022a), we create 63 augmented views per test image. The AdamW optimizer is used with the learning rate set to 0.0005 for the residual parameters in one iteration. Hyperparameters are set as follows: $\alpha=6$, $\beta=5.0$, $\delta=0.95$, $\gamma=0.02$, $\lambda=0.5$, and queue size $M=16$ (since ACE shows resilience to noisy pseudo-labels and benefits from additional samples for class distribution modeling). All experiments are conducted on a single NVIDIA A100 GPU with 80GB of memory.
Results
Robustness to Natural Distribution Shifts.
Our method reliably excels in adapting to varied image datasets with natural shifts, as highlighted in Table 1. Compared to TDA (Karmanov et al. 2024a) and DPE (Zhang et al. 2024d), both our probability and entropy variants outperform them across 4 out of 5 datasets with notable differences, tested across ResNet-50 and ViT-B/16 backbones. Specifically, our ACE achieves top OOD average accuracies of 47.79% with ResNet-50 and 64.63% with ViT-B/16. While DiffTPT (Feng et al. 2023a) occasionally surpasses our method, we found this is due to the limited use of data augmentation in our experiments. Our method also outperforms the few-shot learning technique CoOp (Zhou et al. 2021), underscoring our method’s robust ability to adapt VLMs to OOD scenarios.
Cross-Datasets Generalization.
In Table 2, we evaluate the generality of our ACE-Probability and ACE-Entropy variants across 10 diverse fine-grained datasets, and compare them with state-of-the-art approaches. Due to significant distributional differences among these datasets, performance varies considerably. However, our methods achieve average accuracy improvements of 1.65% and 0.82% over the best-performing methods on the CLIP-ResNet-50 and ViT-B/16 backbones, respectively, surpassing competitors in 6 out of 10 datasets. Notably, on the challenging DTD dataset, characterized by complex texture categories, our approach consistently outperforms prior works across both CLIP backbones by a large margin. Additionally, our approach significantly outperforms the few-shot prompt learning method CoOp (Zhou et al. 2021) in all datasets. These findings highlight the generality and adaptability of our approach in test-time domain transfer, making it highly suitable for real-world applications.
Ablation study
Cache size and zero-shot statistics.
Table 3 presents an ablation study on the impact of zero-shot cache initialization and cache size on the performance of our ACE across five datasets. Table 3 (a) compares DPE with our ACE in two settings: with and without zero-shot (ZS) cache. As shown in the results, ACE with ZS cache achieves higher average accuracies (95.33% on Caltech and 77.30% on Flower with ViT-B/16) compared to DPE, highlighting the benefit of ZS cache initialization. Table 3 (b) examines the effect of cache size (M) using probabilistic (Prob.) and entropy (Ent.) quantities. The results show that larger cache sizes (e.g., M = 16) consistently improve the classification performance of CLIP (95.33% on Caltech and 73.72% on UCF with ViT-B/16). These achievements highlight ACE’s ability in mitigating caching noise and enhancing classification accuracy of VLMs with increased cache capacity.
Setting Strat. Caltech DTD Flower Pets UCF
DPE - 90.83 50.18 67.60 85.97 61.98
ACE w/o ZS cache Prob. 90.55 52.54 69.51 87.41 61.81
ACE w/ ZS cache Prob. 90.91 55.20 70.28 87.95 61.83
ACE w/o ZS cache Ent. 90.39 52.19 69.39 87.00 62.18
ACE w/ ZS cache Ent. 90.99 53.90 70.08 87.54 62.60
DPE - 94.81 54.20 75.07 91.14 70.44
ACE w/o ZS cache Prob. 95.06 56.21 75.80 91.21 72.06
ACE w/ ZS cache Prob. 95.13 56.38 77.47 91.47 73.62
ACE w/o ZS cache Ent. 95.12 56.09 74.99 91.12 72.16
ACE w/ ZS cache Ent. 95.33 56.56 77.30 91.47 73.72
(a) Impact of zero-shot cache
Strat. Cache size Caltech DTD Flower Pets UCF
Prob. M = 3 90.63 51.30 69.96 85.94 60.32
Ent. M = 3 90.91 51.12 68.17 86.02 60.69
Prob. M = 16 90.91 55.20 70.28 87.95 61.83
Ent. M = 16 90.99 53.90 70.08 87.54 62.60
Prob. M = 3 95.09 55.02 77.39 89.94 71.74
Ent. M = 3 95.29 54.31 77.59 90.22 73.35
Prob. M = 16 95.13 56.38 77.47 91.47 73.62
Ent. M = 16 95.33 56.56 77.30 91.47 73.72
(b) Impact of cache size
Table 3: Ablation study on the impact of zero-shot cache initialization and cache size toward ACE.
Visualization of cache features using t-SNE
Figure 3 displays t-SNE (van der Maaten and Hinton 2008) of the cached image features (with cache size $M=6$) for both DPE (Zhang et al. 2024d) and our method on the Food101 (Bossard, Guillaumin, and Van Gool 2014) dataset. Features from 10% of randomly selected classes are shown in different colors, while all others are indicated in gray. The visualizations demonstrate that our priority queue approach successfully gathers high-confidence samples, leading to increasingly representative visual prototypes over time.
Figure 3: t-SNE (van der Maaten and Hinton 2008) visualizations of the stored image features in the cache between DPE (Zhang et al. 2024d) and ACE-Entropy.
Cache reliability
Figure 4 visualizes the cache accuracy and overall classification performance of TDA, DPE, and our ACE across four fine-grained datasets. ACE demonstrates exceptional results, achieving an accuracy of 68.55% on Stanford Cars, surpassing TDA by 1.27% and DPE by 1.24%, and 77.30% on Oxford Flowers, outperforming TDA by 5.88% and DPE by 2.07%. These gains are complemented by ACE’s consistent superiority in cache accuracy across the datasets, reinforcing its ability in enhancing classification power while sustaining reliable and flexible decision boundaries via class-wise thresholding adaptation.
Figure 4: Comparison of cache accuracy vs test accuracy over four datasets between TDA (Karmanov et al. 2024a), DPE (Zhang et al. 2024d), and our proposed ACE-Entropy.
Conclusion
In this paper, we present Adaptive Cache Enhancement (ACE), which decisively modifies class-wise thresholds to ensure the caching of more reliable test samples. Extensive benchmarking across a range of datasets and domains demonstrates that ACE outperforms state-of-the-art methods, particularly in the face of challenging distribution shifts. This approach is not merely practical; it is a robust solution for adapting vision-language models in dynamic environments. Our findings clearly show that effectively addressing noisy cache samples significantly enhances generalizability.
Appendix A Appendix
In appendix, we provide additional details and experimental results to enhance understanding and insights into our proposed method. This supplementary document is organized as follows:
  • $\bullet$
    Detailed Dataset Information: Comprehensive details about the datasets used in our experiments, including their key characteristics and distributions, are provided.
  • $\bullet$
    Textual Prompts Used in Experiments: The text templates used in our experiments for each dataset are listed for reproducibility.
  • $\bullet$
    Full results of ablation study: We presented further details of the ablation study on cache size and ZS cache initialization for all datasets.
  • $\bullet$
    Supplementary for different seeds: We provide the full performance comparison of ACE for 3 different seeds on OOD experiments.
Appendix B Detailed Dataset Information
In Table 4, we provide a comprehensive overview of the datasets employed in our experiments, detailing key statistics such as the number of classes, the sizes of the training, validation, and testing sets, the original tasks for which each dataset was designed. This information enables a thorough understanding of the datasets’ composition and their relevance to the tasks evaluated in our study.
Dataset Classes Training Validation Testing Task
Caltech101 (Fei-Fei, Fergus, and Perona 2007) 100 4,128 1,649 2,465 Object recognition
DTD (Cimpoi et al. 2014) 47 2,820 1,128 1,692 Texture recognition
EuroSAT (Helber et al. 2019) 10 13,500 5,400 8,100 Satellite image recognition
FGVCAircraft (Maji et al. 2013b) 100 3,334 3,333 3,333 Fine-grained aircraft recognition
Flowers102 (Nilsback and Zisserman 2008b) 102 4,093 1,633 2,463 Fine-grained flowers recognition
Food101 (Bossard, Guillaumin, and Van Gool 2014) 101 50,500 20,200 30,300 Fine-grained food recognition
ImageNet (Deng et al. 2009) 1,000 1.28M - 50,000 Object recognition
OxfordPets (Parkhi et al. 2012b) 37 2,944 736 3,669 Fine-grained pets recognition
StanfordCars (Krause et al. 2013b) 196 6,509 1,635 8,041 Fine-grained car recognition
SUN397 (Xiao et al. 2010b) 397 15,880 3,970 19,850 Scene recognition
UCF101 (Soomro, Zamir, and Shah 2012) 101 7,639 1,898 3,783 Action recognition
ImageNet-V2 (Recht et al. 2019) 1,000 - - 10,000 Robustness of collocation
ImageNet-Sketch (Wang et al. 2019) 1,000 - - 50,889 Robustness of sketch domain
ImageNet-A (Hendrycks et al. 2021b) 200 - - 7,500 Robustness of adversarial attack
ImageNet-R (Hendrycks et al. 2021a) 200 - - 30,000 Robustness of multi-domains
Table 4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
Appendix C Textual Prompts Used in Experiments
In Table 5, we present a detailed compilation of the hand-crafted prompts specifically designed and utilized for each dataset in our experiments.
Dataset Prompts
“itap of a {CLASS}.”
ImageNet (Deng et al. 2009) “a bad photo of the {CLASS}.”
ImageNet-V2 (Recht et al. 2019) “a origami {CLASS}.”
ImageNet-Sketch (Wang et al. 2019) “a photo of the large {CLASS}.”
ImageNet-A (Hendrycks et al. 2021b) “a {CLASS} in a video game.”
ImageNet-R (Hendrycks et al. 2021a) “art of the {CLASS}.”
“a photo of the small {CLASS}.”
Caltech101 (Fei-Fei, Fergus, and Perona 2007) “a photo of a {CLASS}.”
DTD (Cimpoi et al. 2014) “{CLASS} texture.”
EuroSAT (Helber et al. 2019) “a centered satellite photo of {CLASS}.”
FGVCAircraft (Maji et al. 2013b) “a photo of a {CLASS}, a type of aircraft.”
Flowers102 (Nilsback and Zisserman 2008b) “a photo of a {CLASS}, a type of flower.”
Food101 (Bossard, Guillaumin, and Van Gool 2014) “a photo of {CLASS}, a type of food.”
OxfordPets (Parkhi et al. 2012b) “a photo of a {CLASS}, a type of pet.”
StanfordCars (Krause et al. 2013b) “a photo of a {CLASS}.”
SUN397 (Xiao et al. 2010b) “a photo of a {CLASS}.”
UCF101 (Soomro, Zamir, and Shah 2012) “a photo of a person doing {CLASS}.”
Table 5: Textual prompts used in experiments. In addition to these prompts, we also employ CuPL (Pratt et al. 2023) prompts to further enhance performance.
Appendix D Full results of ablation study
We report full results of cross-datasets generalization settings to show the effectiveness of zero-shot statistics and choice of cache size in Table 3.
Setting Strat. Aircraft Caltech Cars DTD EuroSAT Flower Food101 Pets SUN397 UCF101 Mean
DPE - 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
w/o ZS cache Prob. 20.97 90.55 59.63 52.54 45.94 69.51 77.47 87.41 64.92 61.81 63.08
w/ ZS cache Prob. 21.33 90.91 59.84 55.20 51.02 70.28 76.83 87.95 65.46 61.83 64.07
w/o ZS cache Ent. 20.37 90.39 59.37 52.19 43.89 69.39 77.48 87.00 64.99 62.18 62.74
w/ ZS cache Ent. 21.48 90.99 60.00 53.90 48.54 70.08 76.91 87.54 65.38 62.60 63.74
DPE - 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
w/o ZS cache Prob. 29.67 95.06 68.44 56.21 55.13 75.80 86.34 91.21 70.40 72.06 70.03
w/ ZS cache Prob. 29.46 95.13 68.72 56.38 54.28 77.47 85.88 91.47 70.83 73.62 70.32
w/o ZS cache Ent. 30.15 95.12 68.40 56.09 54.73 74.99 86.29 91.12 70.51 72.16 69.96
w/ ZS cache Ent. 29.49 95.33 68.55 56.56 53.63 77.30 85.88 91.47 70.88 73.72 70.28
(a) Impact of zero-shot cache
Strat. Cache size Aircraft Caltech Cars DTD EuroSAT Flower Food101 Pets SUN397 UCF101 Mean
Prob. M = 3 20.34 90.63 57.27 51.30 41.02 69.96 76.41 85.94 63.72 60.32 61.69
Ent. M = 3 19.32 90.91 57.08 51.12 42.37 68.17 76.45 86.02 63.33 60.69 61.55
Prob. M = 16 21.33 90.91 59.84 55.20 51.02 70.28 76.83 87.95 65.46 61.83 64.07
Ent. M = 16 21.48 90.99 60.00 53.90 48.54 70.08 76.91 87.54 65.38 62.60 63.74
Prob. M = 3 27.69 95.09 66.26 55.02 54.14 77.39 85.50 89.94 69.38 71.74 69.22
Ent. M = 3 28.68 95.29 66.04 54.31 53.32 77.59 85.50 90.22 69.30 73.35 69.36
Prob. M = 16 29.46 95.13 68.72 56.38 54.28 77.47 85.88 91.47 70.83 73.62 70.32
Ent. M = 16 29.49 95.33 68.55 56.56 53.63 77.30 85.88 91.47 70.88 73.72 70.28
(b) Impact of cache size
Table 6: Ablation study on the impact of zero-shot cache initialization and cache size toward ACE.